# Draw2Cut: Direct On-Material Annotations for CNC Milling

### v1

@@ Abstract.
Figure 1. Left: overview of the three steps for creating artifacts using the Draw2Cut system. Step 1: the user inputs their fabrication intent by drawing on the physical material to be cut. Step 2: the user places the annotated wood in the work space and checks the preview. Step 3: the CNC machine mills the material automatically, allowing users the option of further processing (assemble, paint, etc.). Right: sample artifacts produced with Draw2Cut: a logo, a maze, and a sakura flower.
@@ 1. Introduction
An experienced fabricator (e.g., carpenter) uses tools and experiential techniques to create a scheme to fabricate artifacts. However, this knowledge needs to be gained from experience because it is not codified (tran_oleary_taxon_2021, ). Digital fabrication has attracted considerable attention because of the emergence of personal fabrication (gershenfeld_fab_2005, ; subbaraman_3d_2023, ; mueller_tutorial_2014, ). Numerous computer-aided design (CAD) software have been developed for designing and operating controllable machines for fabrication (bangse_design_2020, ). These tools can bridge the gap between complex fabrication and rapid prototyping (magrisso_digital_2018, ; tran_oleary_taxon_2021, ), enabling designers to convert ideas into objects without the need for handcraft knowledge. However, a significant gap remains between complex fabrication and rapid prototyping (twigg-smith_tools_2021, ; yildirim_digital_2020, ; landwehr_sydow_modding_2022, ), given the limitations identified in the three workflow steps.
@@
The first step involved the measurement and calibration of the materials. Subtractive manufacturing involves using existing materials. Therefore, the size, shape, and relative position of the material in the machine bed before cutting are critical. Typically, CAD users are required to measure or scan materials to extract sufficient data (e.g., length, width, and height) to create a digital canvas as the foundation. Moreover, prior to cutting, CNC users must calibrate the default setting of the machine to align it with the cut target. These tedious tasks are necessary but have little to do with creative design. In this stage, working directly in real space is more intuitive than planning using a CAD interface, because users can quickly obtain a sense of size. Certain projects (e.g., interactive construction (mueller_interactive_2012, )) can be implemented and some commercial tools (e.g., Shaper Trace [^1] https://www.shapertools.com/en-us/trace and Glowforge [^2] https://shop.glowforge.com/products/glowforge-pro) can be operated in real space; however, these approaches require users to draw on a separate interface at a 1:1 scale rather than directly on the material. Working on a separate interface makes 2D planning effortless, but these are limited to subtractive manufacturing products that have a relative spatial relationship with the existing object (e.g.,  (zheng_joinery_2017, )).
@@
The second step is to create a digital blueprint in CAD. The process of producing, modifying, and altering geometries (displayed as vectors) to produce target objects requires the user to master the software. Even experts face challenges when dealing with irregular materials or modifying a blueprint during fabrication (de-los-aires-solis_wood-wood_2023, ), because expanding or modifying the work spaces for machines is difficult (peek_cardboard_2017, ). Instead of editing objects through an interface, it is simpler for users to draw contours in physical space by directly drawing on the material using a pen. Hand-drawing offers users a high degree of freedom to create objects with flexible shapes and sizes. Currently, some projects are focused on a certain product. Maker’s Mark (savage_makers_2015, ) modified a 3D shape for 3D printing using only printouts as primary input. However, this was for additive fabrication; they performed scans to collect processed object spatial information, whereas the subtractive fabrications were required to perform 3D registration to align the virtual space with the real-world coordinates. CopyCAD (follmer_copycad_2010, ) focused on subtractive fabrication by enabling the user to copy and paste real-world objects onto a CNC canvas. However, these methods are limited to two-dimensional (2D) shapes. We explore 3D fabrication methods. Three-axis CNC machines have limited DOF because they can only achieve vertical subtraction. Therefore, we make a trade off between several limitations to investigate limited 3D (extensive x- and y-ranges, limited z-range) subtractive fabrication (rout2022fabrication, ; 2.5D, ).
@@
The third step is toolpath generation. This step can be completed by clicking and selecting options in CAD after finishing the blueprint design, or it can be customized in G-code, which is the most frequently used CNC and 3D printing programming language. This step requires users to be familiar with the cut material and hardware settings (types of milling bits and feed rate) that correspond to the cutting type (e.g., a ball cutter for smooth transitions between cuts, different angle V-bits for V-carving, an end mill cutter for pocketing, and the option to generate straight or arc slots). Previous studies have indicated that this technique is difficult for inexperienced users. Hirsch et al. (hirsch_nothing_2023, ) interviewed 13 professional designers and identified difficulties in computer-aided manufacturing (CAM). For example, to perform an actual cut, enabling only CNC movement followed by the user’s drawing curve, as in a sketch-based CNC (jacobs_sketch-based_2017, ), was insufficient. The machine should be instructed on how to cut (e.g., where to start and how deep to cut). We propose adapting some conventional mark syntax (e.g., cross marks and zigzag marks) from carpenters while making them CNC-readable. This strategy allows us to simplify the interactive process of designing toolpaths while maintaining product diversity.
@@ 2.1. Lowering Technical Barriers for Personal Fabrication
The canonical digital fabrication (twigg-smith_tools_2021, ) comprises several parts. We categorized them into two categories, namely using CAD to design the product in digital form, and using CAM to physically manufacture the output. Typically, CAD is considered an essential component of CAM (what_is_CAM_2021, ). Therefore, focusing on computer-aided design (CAD) for fabrication is a popular research topic. Researchers have added numerous functionalities to software such that users can handle it more easily (peng_–fly_2016, ; peng_d-coil_2015, ). Concurrently, designers have developed CAD interfaces to make visualizations more intuitive and comprehensive (vasquez_jubilee_2020, ; fossdal_interactive_2021, ; moyer2024throwing, ). These studies have focused on software interactions. Another study direction is to remove CAD from the idea-to-fabrication process. Eliminating CAD could decrease the complexity, precision, and reproducibility (peng_roma_2018, ). However, user interactions are not confined to the software. After combining mixed reality, multimodal sensing, and other technologies, various interactive environments or approaches can emerge. This phenomenon lowers the technical barrier for people who want to rapidly prototype creative ideas (that is, personal fabrication). Therefore, we focused on this direction.
Another difficulty encountered after removing the CAD is translating the user’s varied inputs into instructions that the processing machine can execute. CAD provides an interface for users to select the depth, cutting methods, and cutting tools, which are then converted into toolpaths (that is, G-code output files). Consequently, previous CAD studies have focused on toolpath generation. For example, researchers have developed an interface capable of displaying several types of data, not just geometry (tran_oleary_improving_2022, ), allowing for all digital attributes (peng_d-coil_2015, ) and designing toolpaths so that they can be altered parametrically during manufacturing (zoran_human-computer_2013, ). For challenging designs, such as wooden joints, researchers have developed supporting tools, such as interfaces, which can provide a detailed picture of the joints (larsson_tsugite_2020, ) or incorporate a parametric joint library to enable rapid exploration (tian_matchsticks_2018, ). SensiCut incorporates a material-sensing platform that provides a user interface with material information to assist users in determining the appropriate power and speed while avoiding dangerous material cutting (dogan_sensicut_2021, ). If we remove the typical CAD (i.e., screen-free) rather than considering the input method, then we should consider other methods to generate precise G-codes for CNC.
@@ 2.2. Input Methods and Interactive Environment
There are various studies to address the problem of interaction efficiency and alignment accuracy between physical and digital spaces. Previous studies addressed this issue through several approaches. For examples, the multimodal sensing approaches rendered the forms of interactions more diverse. The enriched input types included voice, temperature, pressure, gestures, and touch. FormFab reshapes a thermoplastic material using a heat gun (mueller2019formfab, ), Speaker shapes a wire based on the user’s voice (sensing a simplified sound wave), and Cutter incorporates a thin wire heated by electrical resistance to control the cutter’s pushing or pulling (willis_interactive_2010, ). To adjust the shape of the polyurethane foam beneath the touchscreen, users input the pattern on the touchscreen. Mueller et al. presented an interactive construction (mueller_interactive_2012, ) in which the user uses a laser pointer to draw lines on the screen above the target material within a laser cutter and subsequently cuts the material using a CNC. However, in these studies, rather than sketching directly on a physical item, the user was required to use a separate instrument (such as a drawing pad or heat gun).
@@
The second category includes methods based on augmented reality (AR) (dogan_fabricate_2022, ; dogan_augmented_2024, ). Users physically work on a product using mixed reality (MR) devices as intermediaries to ensure that their instructions are coordinated with the virtual model (iyer_xr-penter_2025, ; arslan_realitycraft_2024, ). For example, SPATA examined integration into virtual design environments using data from two geographical measurements (weichel_spata_2015, ). RoMA allows customers to expand existing items with an AR controller or create well-proportioned physical artifacts (peng_roma_2018, ). AR-based techniques are useful for remote collaboration fabrication (goveia_da_rocha_exquisite_2021, ; kim_machines_2017, ; tian_turn-by-wire_2019, ; tian_adroid_2021, ) and manufacturing, which demand real-time feedback (takahashi_interactive_2016, ; albaugh_augmented_2023, ). For example, in fabrications that require constant adjustments during the process (e.g., clay and ceramics (frost2024sketchpath, )), complex expressions can only be exposed through the embodied process of production (landwehr_sydow_modding_2022, ; tokac_craft-inspired_2022, ). The commercial tool, Shaper Origin [^4] https://www.shapertools.com/en-us/origin, a handheld CNC router, allows a user to cut directly on the material with real-time feedback using an augmented-reality overlay. MR allows continuous or turn-taking inputs, enabling fabrication and design benefits through feedback loops and creative inputs. Because the material we used (e.g., wood) was relatively stable during the process, we did not achieve such real-time interactivity.
@@
However, as stated in the RoMA limitations (peng_roma_2018, ), the fabrication quality and precision of the calibration are difficult to manage because of inaccuracies throughout the overlay procedure between the physical and digital models. Another issue comes from the property in subtractive manufacturing. Adding or deforming fabrication may only require aligning the physical input with the virtual environment and executing computing. The user can deform the material based on the calculated virtual results (mueller2019formfab, ). A 3D printing machine can directly produce outcomes in an organized form (frost2024sketchpath, ). A framework ModelCraft (song2009modelcraft, ) employs a user input approach similar to ours but focuses on physical-to-digital synchronization rather than actual fabrication. However, in subtractive manufacturing, after completing the computation in the virtual space, the system should cut in the real space at the correct point relative to the material to be cut. In other words, the proposed module should perform a secondary calibration to ensure that the virtual model matches the physical space. A project-camera system is required to complete the registration.
@@
For example, CopyCAD (follmer_copycad_2010, ) allows users to copy 2D shapes from real-world objects captured by a camera and project these shapes exactly where the genuine objects are. Maker’s Mark (savage_makers_2015, ) did not incorporate a standard camera, but used 3D scanning to collect spatial information about the processed product. Jacobs et al. developed a sketch-based CNC (jacobs_sketch-based_2017, ), which allows CNC movement, followed by a drawing pattern of the same size as the drawn line. This approach projects an augmented object onto different spaces to enhance interaction techniques. For example, iLamps investigated several interaction strategies to describe object augmentation using a handheld projector (raskar_ilamps_2006, ). Shader Lamps directly recreated the appearance of a real item and boosted its visual quality using a projector (raskar_shader_2001, ). We applied a similar method to the subtractive manufacturing of wood using a camera to compute the environmental information for CNC cutting. In other words, we aligned our physical workspace with the underlying virtual space while calibrating the computed digital results into physical material.
@@ 2.3. Manufacturing Process Execution
Some studies have focused on CAM procedures that simplify the programming process, such as creating a creative coding environment (subbaraman_p5_2022, ). ”Physical-digital programming” first originated for tailoring fabrication workflows for practitioners with varied programming language underpinnings (tran_oleary_physical-digital_2023, ). Based on this approach, Extruder-Turtle was developed to generate G-code for engineers familiar with the open-source Turtle Geometry library (pezutti-dyer_extruder-turtle_2022, ). Researchers developed StickyLand, which organizes code in computational notebooks for data scientists such as Jupyter notebooks (wang_stickyland_2022, ). Their primary objective was to create a programming environment with which users were already comfortable. However, the target users remain constrained to those with programming skills.
Researchers have generalized toolpath editing by inventing Imprimer, which can detect various documents, including interleaved text and pictures, and is suitable for users from various backgrounds (tran_oleary_imprimer_2023, ). This approach is appropriate for patterns that are easily described in literature. However, this approach still struggles with irregular shapes that are difficult to describe verbally. Consequently, sketch-based methods have been devised (bremers_computer_2022, ; willis_spatial_2010, ). SketchPath enables a 3D printer to interpret the user’s hand-drawn textures and use them to organize 3D clay printing shapes (frost2024sketchpath, ). In addition to additive manufacturing, our project obtained a similar method to perform subtractive fabrication using hand-drawn symbols as the cut-type instructions to generate the toolpath.

### v3

@@ Abstract.
Figure 1. Left: Overview of the three steps for creating artifacts using the Draw2Cut system. Step 1: Users input their fabrication intent by drawing on the physical material to be cut. Step 2: Users place the annotated wood in the work space and checks the preview. Step 3: The CNC machine mills the material automatically, allowing users to choose further processing (e.g., assemble, paint). Right: Sample artifacts produced with Draw2Cut: a logo, a maze, and a sakura flower.
@@ 1. Introduction
An experienced fabricator (e.g., carpenter) uses tools and experiential techniques to create a scheme to fabricate artifacts. However, this knowledge needs to be gained from experience because it is not codified (Tran O’Leary et al., 2021). Digital fabrication has attracted considerable attention because of the emergence of personal fabrication (Gershenfeld, 2005; Subbaraman and Peek, 2023; Mueller et al., 2014). Numerous computer-aided design (CAD) software have been developed for designing and operating controllable machines for fabrication (Bangse et al., 2020). These tools can bridge the gap between complex fabrication and rapid prototyping (Magrisso et al., 2018; Tran O’Leary et al., 2021), enabling designers to convert ideas into objects without the need for handcraft knowledge. However, a significant gap remains between complex fabrication and rapid prototyping (Twigg-Smith et al., 2021; Yildirim et al., 2020; Landwehr Sydow et al., 2022), given the limitations identified in the three workflow steps.
@@
The first step involved the measurement and calibration of the materials. Subtractive manufacturing involves using existing materials. Therefore, the size, shape, and relative position of the material in the machine bed before cutting are critical. Typically, CAD users are required to measure or scan materials to extract sufficient data (e.g., length, width, and height) to create a digital canvas as the foundation. Moreover, prior to cutting, CNC users must calibrate the default setting of the machine to align it with the cut target. These tedious tasks are necessary but have little to do with creative design. In this stage, working directly in real space is more intuitive than planning using a CAD interface, because users can quickly obtain a sense of size. Certain projects (e.g., interactive construction (Mueller et al., 2012)) can be implemented and some commercial tools (e.g., Shaper Trace [^1] https://www.shapertools.com/en-us/trace and Glowforge [^2] https://shop.glowforge.com/products/glowforge-pro) can be operated in real space; however, these approaches require users to draw on a separate interface at a 1:1 scale rather than directly on the material. Working on a separate interface makes 2D planning effortless, but these are limited to subtractive manufacturing products that have a relative spatial relationship with the existing object (e.g.,  (Zheng et al., 2017)).
@@
The second step is to create a digital blueprint in CAD. The process of producing, modifying, and altering geometries (displayed as vectors) to produce target objects requires the user to master the software. Even experts face challenges when dealing with irregular materials or modifying a blueprint during fabrication (de-los Aires-Solís and Gonzalez-Quintial, 2023), because expanding or modifying the work spaces for machines is difficult (Peek et al., 2017). Instead of editing objects through an interface, it is simpler for users to draw contours in physical space by directly drawing on the material using a pen. Hand-drawing offers users a high degree of freedom to create objects with flexible shapes and sizes. Currently, some projects are focused on a certain product. Maker’s Mark (Savage et al., 2015) modified a 3D shape for 3D printing using only printouts as primary input. However, this was for additive fabrication; they performed scans to collect processed object spatial information, whereas the subtractive fabrications were required to perform 3D registration to align the virtual space with the real-world coordinates. CopyCAD (Follmer et al., 2010) focused on subtractive fabrication by enabling the user to copy and paste real-world objects onto a CNC canvas. However, these methods are limited to two-dimensional (2D) shapes. We explore 3D fabrication methods. Three-axis CNC machines have limited DOF because they can only achieve vertical subtraction. Therefore, we make a trade off between several limitations to investigate limited 3D (extensive x- and y-ranges, limited z-range) subtractive fabrication (Rout and Patel, 2022; CNC, 2024).
@@
The third step is toolpath generation. This step can be completed by clicking and selecting options in CAD after finishing the blueprint design, or it can be customized in G-code, which is the most frequently used CNC and 3D printing programming language. This step requires users to be familiar with the cut material and hardware settings (types of milling bits and feed rate) that correspond to the cutting type (e.g., a ball cutter for smooth transitions between cuts, different angle V-bits for V-carving, an end mill cutter for pocketing, and the option to generate straight or arc slots). Previous studies have indicated that this technique is difficult for inexperienced users. Hirsch et al. (Hirsch et al., 2023) interviewed 13 professional designers and identified difficulties in computer-aided manufacturing (CAM). For example, to perform an actual cut, enabling only CNC movement followed by the user’s drawing curve, as in a sketch-based CNC (Jacobs, 2017), was insufficient. The machine should be instructed on how to cut (e.g., where to start and how deep to cut). We propose adapting some conventional mark syntax (e.g., cross marks and zigzag marks) from carpenters while making them CNC-readable. This strategy allows us to simplify the interactive process of designing toolpaths while maintaining product diversity.
@@ 2.1. Lowering Technical Barriers for Personal Fabrication
The canonical digital fabrication (Twigg-Smith et al., 2021) comprises several parts. We categorized them into two categories, namely using CAD to design the product in digital form, and using CAM to physically manufacture the output. Typically, CAD is considered an essential component of CAM (Deans, 2021). Therefore, focusing on computer-aided design (CAD) for fabrication is a popular research topic. Researchers have added numerous functionalities to software such that users can handle it more easily (Peng et al., 2016, 2015). Concurrently, designers have developed CAD interfaces to make visualizations more intuitive and comprehensive (Vasquez et al., 2020; Fossdal et al., 2021; Moyer et al., 2024). These studies have focused on software interactions. Another study direction is to remove CAD from the idea-to-fabrication process. Eliminating CAD could decrease the complexity, precision, and reproducibility (Peng et al., 2018). However, user interactions are not confined to the software. After combining mixed reality, multimodal sensing, and other technologies, various interactive environments or approaches can emerge. This phenomenon lowers the technical barrier for people who want to rapidly prototype creative ideas (that is, personal fabrication). Therefore, we focused on this direction.
Another difficulty encountered after removing the CAD is translating the user’s varied inputs into instructions that the processing machine can execute. CAD provides an interface for users to select the depth, cutting methods, and cutting tools, which are then converted into toolpaths (that is, G-code output files). Consequently, previous CAD studies have focused on toolpath generation. For example, researchers have developed an interface capable of displaying several types of data, not just geometry (Tran O’Leary et al., 2022), allowing for all digital attributes (Peng et al., 2015) and designing toolpaths so that they can be altered parametrically during manufacturing (Zoran et al., 2013). For challenging designs, such as wooden joints, researchers have developed supporting tools, such as interfaces, which can provide a detailed picture of the joints (Larsson et al., 2020) or incorporate a parametric joint library to enable rapid exploration (Tian et al., 2018). SensiCut incorporates a material-sensing platform that provides a user interface with material information to assist users in determining the appropriate power and speed while avoiding dangerous material cutting (Dogan et al., 2021). If we remove the typical CAD (i.e., screen-free) rather than considering the input method, then we should consider other methods to generate precise G-codes for CNC.
@@ 2.2. Input Methods and Interactive Environment
There are various studies to address the problem of interaction efficiency and alignment accuracy between physical and digital spaces. Previous studies addressed this issue through several approaches. For examples, the multimodal sensing approaches rendered the forms of interactions more diverse. The enriched input types included voice, temperature, pressure, gestures, and touch. FormFab reshapes a thermoplastic material using a heat gun (Mueller et al., 2019), Speaker shapes a wire based on the user’s voice (sensing a simplified sound wave), and Cutter incorporates a thin wire heated by electrical resistance to control the cutter’s pushing or pulling (Willis et al., 2010b). To adjust the shape of the polyurethane foam beneath the touchscreen, users input the pattern on the touchscreen. Mueller et al. presented an interactive construction (Mueller et al., 2012) in which the user uses a laser pointer to draw lines on the screen above the target material within a laser cutter and subsequently cuts the material using a CNC. However, in these studies, rather than sketching directly on a physical item, the user was required to use a separate instrument (such as a drawing pad or heat gun).
@@
The second category includes methods based on augmented reality (AR) (Dogan et al., 2022a, 2024). Users physically work on a product using mixed reality (MR) devices as intermediaries to ensure that their instructions are coordinated with the virtual model (Iyer et al., 2025; Arslan et al., 2024). For example, SPATA examined integration into virtual design environments using data from two geographical measurements (Weichel et al., 2015). RoMA allows customers to expand existing items with an AR controller or create well-proportioned physical artifacts (Peng et al., 2018). AR-based techniques are useful for remote collaboration fabrication (Goveia da Rocha et al., 2021; Kim et al., 2017; Tian et al., 2019; Tian and Paulos, 2021) and manufacturing, which demand real-time feedback (Takahashi and Miyashita, 2016; Albaugh et al., 2023). For example, in fabrications that require constant adjustments during the process (e.g., clay and ceramics (Frost et al., 2024)), complex expressions can only be exposed through the embodied process of production (Landwehr Sydow et al., 2022; Tokac et al., 2022). The commercial tool, Shaper Origin [^4] https://www.shapertools.com/en-us/origin, a handheld CNC router, allows a user to cut directly on the material with real-time feedback using an augmented-reality overlay. MR allows continuous or turn-taking inputs, enabling fabrication and design benefits through feedback loops and creative inputs. Because the material we used (e.g., wood) was relatively stable during the process, we did not achieve such real-time interactivity.
@@
However, as stated in the RoMA limitations (Peng et al., 2018), the fabrication quality and precision of the calibration are difficult to manage because of inaccuracies throughout the overlay procedure between the physical and digital models. Another issue comes from the property in subtractive manufacturing. Adding or deforming fabrication may only require aligning the physical input with the virtual environment and executing computing. The user can deform the material based on the calculated virtual results (Mueller et al., 2019). A 3D printing machine can directly produce outcomes in an organized form (Frost et al., 2024). A framework ModelCraft (Song et al., 2009) employs a user input approach similar to ours but focuses on physical-to-digital synchronization rather than actual fabrication. However, in subtractive manufacturing, after completing the computation in the virtual space, the system should cut in the real space at the correct point relative to the material to be cut. In other words, the proposed module should perform a secondary calibration to ensure that the virtual model matches the physical space. A project-camera system is required to complete the registration.
@@
For example, CopyCAD (Follmer et al., 2010) allows users to copy 2D shapes from real-world objects captured by a camera and project these shapes exactly where the genuine objects are. Maker’s Mark (Savage et al., 2015) did not incorporate a standard camera, but used 3D scanning to collect spatial information about the processed product. Jacobs et al. developed a sketch-based CNC (Jacobs, 2017), which allows CNC movement, followed by a drawing pattern of the same size as the drawn line. This approach projects an augmented object onto different spaces to enhance interaction techniques. For example, iLamps investigated several interaction strategies to describe object augmentation using a handheld projector (Raskar et al., 2006). Shader Lamps directly recreated the appearance of a real item and boosted its visual quality using a projector (Raskar et al., 2001). We applied a similar method to the subtractive manufacturing of wood using a camera to compute the environmental information for CNC cutting. In other words, we aligned our physical workspace with the underlying virtual space while calibrating the computed digital results into physical material.
@@ 2.3. Manufacturing Process Execution
Some studies have focused on CAM procedures that simplify the programming process, such as creating a creative coding environment (Subbaraman and Peek, 2022). ”Physical-digital programming” first originated for tailoring fabrication workflows for practitioners with varied programming language underpinnings (Tran O’Leary et al., 2023b). Based on this approach, Extruder-Turtle was developed to generate G-code for engineers familiar with the open-source Turtle Geometry library (Pezutti-Dyer and Buechley, 2022). Researchers developed StickyLand, which organizes code in computational notebooks for data scientists such as Jupyter notebooks (Wang et al., 2022). Their primary objective was to create a programming environment with which users were already comfortable. However, the target users remain constrained to those with programming skills.
Researchers have generalized toolpath editing by inventing Imprimer, which can detect various documents, including interleaved text and pictures, and is suitable for users from various backgrounds (Tran O’Leary et al., 2023a). This approach is appropriate for patterns that are easily described in literature. However, this approach still struggles with irregular shapes that are difficult to describe verbally. Consequently, sketch-based methods have been devised (Bremers, 2022; Willis et al., 2010a). SketchPath enables a 3D printer to interpret the user’s hand-drawn textures and use them to organize 3D clay printing shapes (Frost et al., 2024). In addition to additive manufacturing, our project obtained a similar method to perform subtractive fabrication using hand-drawn symbols as the cut-type instructions to generate the toolpath.